# As Soon as Possible but Rationally

### v4

@@ 1 Introduction
PO, weights $\Pi_{2}^{\mathsf{P}}$-complete $\mathsf{PSPACE}$-complete $\mathsf{PSPACE}$-complete $\mathsf{NEXPTIME}$-complete [gaspard-pareto-quanti-reach]
@@
NE, weights $\mathsf{coNP}$-complete $\mathsf{coNP}$-complete $\mathsf{NP}$-complete Unknown, $\mathsf{EXPTIME}$-hard1
@@
The results for PO rationality are as follows. (1) For the verification problems, we assume that the behavior of the system is formalized by a nondeterministic Mealy machine, used to represent a (usually infinite) set of its possible implementations. For each of those implementations, we verify that the quantitative reachability objective of the system is met against any rational behavior of the environment. We establish that this problem is $\mathsf{PSPACE}$-complete. To obtain the upper bound, we rely on a genuine combination of techniques based on Parikh automata and a recursive $\mathsf{PSPACE}$ algorithm (for positive Boolean combinations of bounded safety objectives, a problem of independent interest). Parikh automata are used to guess a compact representation of certificates which are paths of possibly exponential length in the size of the problem input. When the Mealy machine is deterministic, we show that the complexity goes down to $\Pi_{2}^{\mathsf{P}}$-completeness, as the previous $\mathsf{PSPACE}$ algorithm is replaced by a $\mathsf{coNP}$ oracle. (2) For the synthesis problems, we only consider the cooperative version which we prove to be $\mathsf{PSPACE}$-complete, as the non-cooperative version was established to be $\mathsf{NEXPTIME}$-complete in [gaspard-pareto-quanti-reach].
@@ 3 Studied Problems
Our main results for Problems 1-2 are the following ones when the rational responses of the environment are 00-fixed PO plays. One problem was already solved in [gaspard-pareto-quanti-reach].
@@
  4. (d)
    The Non-Cooperative Pareto Synthesis problem is $\mathsf{NEXPTIME}$-complete [gaspard-pareto-quanti-reach].

### v6

@@ 1 Introduction
PO, weights $\Pi_{2}^{\mathsf{P}}$-complete $\mathsf{PSPACE}$-complete $\mathsf{PSPACE}$-complete Unknown, $\mathsf{NEXPTIME}$-hard1  [gaspard-pareto-quanti-reach]
@@
NE, weights $\mathsf{coNP}$-complete $\mathsf{coNP}$-complete $\mathsf{NP}$-complete Unknown, $\mathsf{EXPTIME}$-hard2
@@
    In [gaspard-pareto-quanti-reach], the authors show the $\mathsf{NEXPTIME}$-completeness when the weight function is the same for each player, we do not have the result for the general case of multiple weight functions.
  • 2
@@
The results for PO rationality are as follows. (1) For the verification problems, we assume that the behavior of the system is formalized by a nondeterministic Mealy machine, used to represent a (usually infinite) set of its possible implementations. For each of those implementations, we verify that the quantitative reachability objective of the system is met against any rational behavior of the environment. We establish that this problem is $\mathsf{PSPACE}$-complete. To obtain the upper bound, we rely on a genuine combination of techniques based on Parikh automata and a recursive $\mathsf{PSPACE}$ algorithm (for positive Boolean combinations of bounded safety objectives, a problem of independent interest). Parikh automata are used to guess a compact representation of certificates which are paths of possibly exponential length in the size of the problem input. When the Mealy machine is deterministic, we show that the complexity goes down to $\Pi_{2}^{\mathsf{P}}$-completeness, as the previous $\mathsf{PSPACE}$ algorithm is replaced by a $\mathsf{coNP}$ oracle. (2) For the synthesis problems, we only consider the cooperative version which we prove to be $\mathsf{PSPACE}$-complete, as the non-cooperative version was established to be $\mathsf{NEXPTIME}$-complete in [gaspard-pareto-quanti-reach] when there is a common weight function for each player.
@@ 3 Studied Problems
Our main results for Problems 3-3 are the following ones when the rational responses of the environment are 00-fixed PO plays. One problem, the Non-Cooperative Pareto Synthesis problem, was already partially solved in [gaspard-pareto-quanti-reach] [^3] In [gaspard-pareto-quanti-reach], the authors show the $\mathsf{NEXPTIME}$-completeness when there is one single common weight function for all players. We have no result on an $\mathsf{NEXPTIME}$ membership for the general case so far..